John’s Weed-B-Gone yard service sprays lawns to rid them of weeds. John mixes the two chemicals, Weed-X and Pest-O, in proportions depending on the climate and the particular weed problems of the season. A standard mix calls for a gallon of the mixture to combine equal parts of Weed-X and Pest-O. Weed-X has a standard cost of $10.40 per gallon and Pesto-O has a standard cost of $25.20 per gallon. Each gallon can treat 100 square yards of lawn. John expects to treat 535,000 square yards of lawn.

 

For the past season, John treated 646,000 square yards. He used 3,600 gallons of Weed-X and 3,400 gallons of Pest-O. He paid a total of $30,960 for the Weed-X and $82,960 for the Pest-O.

 

Required:

a. Compute the materials price and efficiency variances for the season.

b. Compute the materials mix and yield variances for the season.
Answer #1

step 1)  (1/100)/2= 0.005

30,960/ 3600= 8.6 or 8.60

82,960/ 3,400= 24.4 or 24.40


3,600+3,400= 7,000

Weed X

8.60*3,600= 30,960

10.40* 3,600= 37,440

10.40* (1/2* 7,000)= 36,400

10.40* (0.005* 646,000)= 33,592


pesto

24.40* 3,400= 82,960

25.20* 3,400= 85,680

25.20* (1/2* 7,000)= 88,200

25.20* (0.005* 646,000)= 81,396


total

30,960+ 82,960= 113,920

37,440+ 85,680= 123,120

36,400+ 88,200= 124,600

33,592+ 81,396= 114,988


part A) materials price    Answer:

weed x= 37,440- 30,960= 6,480 F

pest-o= 85,680- 82,960= 2,720 F

total 6,480+ 2,720= 9,200 F


Efficiency variance   Answer

weed X: 33,593- 37,440= 3,848 U

Pest-O: 81,396- 85,680= 4,284 U

total: 3,848+ 4,284= 8,132 U


Part B) Materials mix

weed X: 36,400- 37,440= 1,040 U

Pest-O: 88,200- 85,680= 2,520 F

total: 1,040- 2,520= 1,480 F 


Yield variances

weed X: 33,592- 36,400= 2,808 U

pest-o: 81,396- 88,200= 6,804 U

total: 2,808+ 6,804= 9,612 U